# Multi-scale Graph Autoregressive Modeling: Molecular Property Prediction via Next Token Prediction

### v1

@@ 1 Introduction
Molecular property prediction is a core challenge in drug discovery (wu2018moleculenet) and has increasingly become a primary benchmark for “Foundation Models (FM) for Science.” (khan2025comprehensive) Within FM paradigm, the training recipe has standardized: decoder-only Transformers trained via next-token prediction (NTP) serve as the dominant engine for scale and generalization (wang2024emu3; xia2025nature; zhang2025unigenx). Consequently, as the core components of these training stacks become heavily optimized, the pragmatic research focus shifts away from bespoke backbone re-design toward optimizing the input representation to fit this proven architecture (touvron2023llama). However, a critical gap remains: current NTP-based molecular models, which largely rely on 1D SMILES strings, fail to explicitly capture graph topology and consequently trail behind specialist graph models in predictive accuracy (xia2025nature).
@@
Conversely, graph-native methods explicitly model topology but typically depart from pure decoder-only NTP, instead adopting hybrid architectures or objectives that involve, to varying degrees, input corruption (e.g., masked node prediction) (li2023knowledge; shehzad2024graph; lu2025uni; kong2025unimomo). While sophisticated strategies—such as incorporating global knowledge nodes or avoiding random masking—have been developed to mitigate the information loss caused by masking (li2023knowledge; rong2020self; liu2024mask; you2020graph; xu2021self), these methods remain fundamentally bound by the corruption-coverage trade-off (wettig2023should). This limitation is particularly acute in activity-cliff scenarios, where a single atomic change triggers a drastic property shift; masking the local neighborhood of a critical functional group effectively removes the precise evidence needed to discern such subtle differences. Standard NTP, by contrast, offers a dense supervision signal without corrupting the visible prefix context (clark2020electra), yet it lacks a representation interface that exposes graph connectivity as effectively as text.
@@
To bridge this gap, we propose Connection-Aware Motif Sequencing (CamS), a tokenizer-level interface that makes molecular graphs directly learnable by standard decoder-only NTP through a three-stage design. First, to ensure information efficiency while preserving semantics, we adapt byte-pair encoding (BPE)-style mining on molecular graphs (shibata1999byte; geng2023novo; shen2024graphbpe). This produces motif vocabulary with frequent merge operation and introduces a tunable molecular scale that allows direct granularity control. To ensure faithful encoding of fine-grained chemical detail, we further integrate a Single-Atom Vocabulary Closure (SAVC) mechanism. Second, to make the graph autoregressive-ready, we partition the molecule into a BPEGraph at a specific scale and serialize it via scaffold-rooted breadth-first search (BFS)  (bundy1984breadth). This produces a CamS subsequence with a stable Intra-scale Order, moving from the global core to peripheral functional groups. Third, and most critically, to construct a hierarchical context, we concatenate subsequences from fine to coarse. This strategy circumvents the trade-off between scales and establishes an Inter-scale Order in the final CamS sequence, enabling high-level motifs to be predicted by conditioning on dense, uncorrupted fine-scale evidence. Collectively, these steps unify substructure compression and serialization through a dual-causal ordering—sequencing motifs within each scale and stacking scales hierarchically—to enable the model to comprehend global scaffolds based on explicit, uncorrupted local structural evidence.
@@
Built on CamS, we instantiate CamS-LLaMA by pre-training a native decoder-only backbone with standard NTP on CamS sequences, keeping the architecture and objective unchanged (touvron2023llama). To keep attribution controlled, we inject only a lightweight molecular fingerprint (cereto2015molecular) prior during downstream fine-tuning, while keeping pre-training purely sequence/structure-driven. Across MoleculeNet (wu2018moleculenet) and the activity-cliff stress test MoleculeACE (van2022exposing), CamS-LLaMA achieves state-of-the-art performance at a comparable model scale to strong graph self-supervised learning baselines, despite relying on substantially weaker priors (li2023knowledge). Beyond accuracy, we provide mechanism-level evidence. Targeted ablation studies confirm that multi-scale concatenation is the key driver of performance, while interpretability analysis reveals that the causal serialization effectively drives attention toward cliff-determining structural differences.
@@ 2.1 Molecular Property Prediction within Foundation Models Paradigm
Molecular property prediction is increasingly adopting a foundation model approach: large-scale Transformer pre-training followed by fine-tuning for diverse property endpoints (awais2025foundation; xia2025nature). One approach adopts standard large language model (LLM) methodology by representing molecules as SMILES strings and pre-training decoder-only models via NTP (xia2025nature; cai2025chemfm). This leverages mature LLM infrastructure and enables straightforward large-scale training. However, SMILES is not structure-native, weakening topology-based signals. String-based edits often misalign with actual chemical changes, and performance consistently trails strong graph-native methods on property prediction tasks (xia2025nature). Alternatively, graph-native foundation models directly process molecular graphs to preserve connectivity and local topology (you2021graph; liupre; you2020graph; xu2021self). High-performing methods typically employ encoder-only Graph Transformers (shehzad2024graph) designed for property prediction. While effective, these models deviate from the vanilla decoder-only NTP approach that exhibits superior generalization and scalability (li2023knowledge; rong2020self). More recent structural foundation models integrate graphs with autoregressive components but frequently depend on additional objectives (often incorporating 3D structure) and learned molecular encoders rather than pure tokenizer-level NTP (kong2025unimomo; lu2025uni; zhang2025unigenx).
@@ 2.2 Molecular Tokenization Strategies
Within a foundation model paradigm, tokenizer design becomes a key interface for scaling LLMs. In NLP, BPE yields strong compression via simple frequency-based merges without training an extra tokenizer model. Molecular tokenization spans string-based and graph-based methods (wang2019smiles; liu2023rethinking). String-based schemes operate on SMILES (weininger1988smiles) or related linearizations (krenn2022selfies), while graph-based schemes have evolved from basic node/edge-level units to fragment-level motifs (liu2023rethinking). Some rule-based motif tokenizations rely on limited handcrafted chemistry rules (zhang2021motif); triplet tokenizations, as used in KPGT (Knowledge-guided Pre-training of Graph Transformer), show that enriching a node token with only one extra atom can substantially boost property prediction (li2023knowledge); ring-or-path-based (wollschlager2024expressivity) and tree-based (jin2018junction) schemes also expose substructures at different granularities; some other works learn tokenization via trainable rules (sun2024representing) or by training a GNN as a tokenizer (liu2023rethinking; luonode). Particularly noteworthy are data-driven motif mining approaches that adapt BPE idea to graphs to obtain reusable, multi-scale fragments with strong compression and minimal reliance on manually designed rules or additional trained models (geng2023novo; shen2024graphbpe; kong2022molecule). In addition, cross-level tokenization schemes, which combine motifs at different structural levels, also provide complementary design ideas (ji2022relmole; chenhierarchical).
@@ 3 Method
Our molecular representation must satisfy two requirements: (1) chemical fidelity, i.e., sensitivity to pharmacophore-level structures and their topology; and (2) autoregressive foundation model compatibility, i.e., a compressed interface naturally aligned with an NTP paradigm (touvron2023llama). Motivated by this, we propose Connection-Aware Motif Sequencing (CamS), which segments molecular graphs into reusable frequent motifs and, crucially, serializes the motif graph into multi-scale connection-aware causal sequences compatible with NTP pre-training (Section 3.1). Built on CamS, CamS-LLaMA performs large-scale autoregressive pre-training and injects a only lightweight fingerprint prior during downstream fine-tuning (Section 3.2). Finally, we analyze the characteristics and advantages of multi-scale CamS autoregressive modeling by comparing it with Graphformer-style modeling (Section 3.3).
@@ 3.1.1 Substructure Level: CamS Vocabulary
BPE-style Mining. Following (geng2023novo; shen2024graphbpe), we first learn an ordered list of merge operations $\mathcal{O}$ from co-occurrence statistics. We then apply $\mathcal{O}$ to the corpus to mine the resulting subgraphs. This yields the BPEMotif set (frequent subgraphs with connectivity patterns) and the base BPE-SAV (Single-Atom Vocabulary with connectivity patterns) naturally captured during mining.
@@ 3.1.2 Per-Scale Encoding
Serialization (Intra-scale Order). To adapt the graph for decoder-only NTP, we serialize $V_{s}$ into an ordered list via Breadth-First Search (BFS). We select the motif with the largest atom count (typically the core scaffold) as the root. This strategy prioritizes global backbone structure before local substitutions (zhang2021motif), establishing a stable Center-to-Periphery causal order.
@@ 3.2.1 Autoregressive Pre-training via NTP
CamS-LLaMA adopts a standard LLaMA-style decoder (touvron2023llama); in pre-training, every CamS token list is treated as an ordinary token sequence in the shared vocabulary $\Sigma$ and is consumed by the backbone under the usual causal mask, with no graph-specific architectural modification. Concretely, each molecule yields five CamS sequence views $\{X_{1},X_{7},X_{27},X_{67},X\}$: four single-scale BFS sequences tokenized with $\Sigma_{1\mathrm{K}}/\Sigma_{7\mathrm{K}}/\Sigma_{27\mathrm{K}}/\Sigma_{67\mathrm{K}}$, plus the concatenated multi-scale sequence $X$ in Eq. (1); we treat each view as one NTP training instance (details in Appendix A.3). Given any view $X=(x_{1},\ldots,x_{T})$, token embeddings form $\mathbf{H}^{(0)}$, which is transformed by $L$ causal self-attention blocks into $\mathbf{H}^{(L)}$; a tied output projection then yields $p_{\theta}(x_{t+1}\mid x_{\leq t})$.
@@ 3.2.2 Fine-tuning with Fingerprint Injection
Many molecular pre-training methods inject rich handcrafted descriptors early (e.g., during pre-training) (li2023knowledge). In contrast, we keep CamS as the primary representation and inject only one widely-used structural prior–a fingerprint–only during fine-tuning, to make the contribution of CamS representations more attributable and controlled.
@@
Graphformer (Flat/Static Bias). $\mathcal{V}_{\text{tok}}$ corresponds to single-scale atoms or triplets. $\psi_{ij}$ encodes static structural biases (e.g., shortest-path-distance bias injected into attention) independent of token content, yielding an isotropic receptive field at a given graph distance (ying2021transformers).
@@
Sketch. The inequality follows from the Data Processing Inequality (DPI): masking is a stochastic post-processing of $Z_{t}$ that cannot increase mutual information (cover1999elements). In graph MNP, this loss of information is exacerbated by evidence-pattern uncertainty: the most relevant local-neighborhood evidence for a masked token may be co-masked, making the conditional distribution unstable across masking patterns. Recent graph pretraining works explicitly mitigate this by masking structured units (e.g., triplets / local subgraphs) or designing structure-guided masking, and by injecting global knowledge tokens (li2023knowledge; rong2020self; liu2024mask). In contrast, CamS NTP for coarse-scale tokens conditions on an uncorrupted fine-to-coarse history, avoiding such stochastic evidence loss. This perspective is consistent with classic critiques of Masked Language Modelling (MLM): due to input corruption and factorized prediction over masked positions, it introduces a pre-train–fine-tune discrepancy and neglects dependencies among masked targets (yang2019xlnet).
@@
This trade-off (more predicted tokens vs. more corruption) is central to masked modeling (wettig2023should). Empirically, KPGT finds the best masking ratio can be as high as $\rho=0.5$ when using triplet-style tokens and global knowledge nodes to stabilize context (li2023knowledge). Meanwhile, dense token-level learning signals are known to substantially improve sample efficiency compared to sparse masked-only supervision (clark2020electra). CamS further strengthens this effect by providing multi-view supervision: the same molecule contributes token-level targets repeatedly across scales.
@@ 4.1 Experimental Setup
CamS-Tokenizer Training. CamS-tokenizer training is expensive because replaying $\mathcal{O}$ to materialize $\mathcal{V}$ repeatedly invokes RDKit’s subgraph-isomorphism matching (the dominant bottleneck) (landrum2021rdkit; ehrlich2011maximum) to extract motifs and attachment patterns, which is difficult to accelerate. We therefore train the tokenizer once on ChEMBL-34 (gaulton2017chembl), a moderately sized yet chemically diverse corpus, at the largest scale (67K). Concretely, we mine an ordered merge list $\mathcal{O}=(o_{1},\dots,o_{K})$ with K=685K{=}685 merges, apply it to obtain $\mathcal{V}$, and then perform SAVC to ensure single-atom coverage for common elements. Smaller scales are obtained without retraining by prefix slicing: each scale uses a prefix $\mathcal{O}_{\leq k}$ together with its induced vocabulary subset (i.e., motifs mined by applying $\mathcal{O}_{\leq k}$), where 1K uses k=0k{=}0, 7K/27K use k=62/210k{=}62/210, and 67K uses k=685k{=}685. Implementation details and full procedures are provided in Appendix A.1.
Pre-training Dataset and Configuration. We instantiate a 16-layer LLaMA-style decoder (hidden size 720, 8 attention heads) using HuggingFace Transformers. For self-supervised pre-training, we use Enamine675M (675M molecules) (shivanyuk2007enamine) and optimize NTP (Section 3.2.1). For each molecule, we construct five CamS training sequences: four single-scale BFS sequences tokenized with 1K/7K/27K/67K vocabularies, plus the concatenated multi-scale sequence in Eq. (1). We treat each sequence as one training instance and uniformly shuffle all instances, resulting in roughly $5\times$ as many sequences as molecules. All optimization details (schedule, batch size, steps) are summarized in Appendix C.1.
Downstream Benchmarks and Baselines. We evaluate on MoleculeNet (11 tasks; broad endpoints that probe general-purpose molecular property prediction) (wu2018moleculenet) and MoleculeACE (30 activity-cliff tasks; stress-testing fine-grained structural sensitivity under subtle edits that induce large activity shifts) (van2022exposing). To enable an apples-to-apples comparison against a recent strong pre-training baseline, we adopt the data splits and evaluation protocol used in KPGT (li2023knowledge). On MoleculeNet, following KPGT’s setup, we compare primarily against self-supervised pre-training baselines to isolate the effect of (1) representation/tokenization and (2) the pre-training objective. On MoleculeACE, we report the benchmark’s comprehensive ML/DL baselines, since activity cliffs emphasize discriminating small structural changes where strong feature-based methods remain competitive. Unless stated otherwise, CamS-LLaMA is kept at $\sim$100M parameters (comparable to KPGT); fingerprints are injected only during fine-tuning as a minimal prior, while pre-training is purely sequence-based.

### v2

@@ 1 Introduction
Molecular property prediction is a core challenge in drug discovery (Wu et al., 2018) and has increasingly become a primary benchmark for “Foundation Models (FM) for Science.” (Khan et al., 2025) Within FM paradigm, the training recipe has standardized: decoder-only Transformers trained via next-token prediction (NTP) serve as the dominant engine for scale and generalization (Wang et al., 2024; Xia et al., 2025; Zhang et al., 2025). Consequently, as the core components of these training stacks become heavily optimized, the pragmatic research focus shifts away from bespoke backbone re-design toward optimizing the input representation to fit this proven architecture (Touvron et al., 2023). However, a critical gap remains: current NTP-based molecular models, which largely rely on 1D SMILES strings, fail to explicitly capture graph topology and consequently trail behind specialist graph models in predictive accuracy (Xia et al., 2025).
@@
Conversely, graph-native methods explicitly model topology but typically depart from pure decoder-only NTP, instead adopting hybrid architectures or objectives that involve, to varying degrees, input corruption (e.g., masked node prediction) (Li et al., 2023; Shehzad et al., 2024; Lu et al., 2025; Kong et al., 2025). While sophisticated strategies—such as incorporating global knowledge nodes or avoiding random masking—have been developed to mitigate the information loss caused by masking (Li et al., 2023; Rong et al., 2020; Liu et al., 2024; You et al., 2020; Xu et al., 2021), these methods remain fundamentally bound by the corruption-coverage trade-off (Wettig et al., 2023). This limitation is particularly acute in activity-cliff scenarios, where a single atomic change triggers a drastic property shift; masking the local neighborhood of a critical functional group effectively removes the precise evidence needed to discern such subtle differences. Standard NTP, by contrast, offers a dense supervision signal without corrupting the visible prefix context (Clark et al., 2020), yet it lacks a representation interface that exposes graph connectivity as effectively as text.
@@
To bridge this gap, we propose Connection-Aware Motif Sequencing (CamS), a tokenizer-level interface that makes molecular graphs directly learnable by standard decoder-only NTP through a three-stage design. First, to ensure information efficiency while preserving semantics, we adapt byte-pair encoding (BPE)-style mining on molecular graphs (Shibata et al., 1999; Geng et al., 2023; Shen and Poczos, ). This produces motif vocabulary with frequent merge operation and introduces a tunable molecular scale that allows direct granularity control. To ensure faithful encoding of fine-grained chemical detail, we further integrate a Single-Atom Vocabulary Closure (SAVC) mechanism. Second, to make the graph autoregressive-ready, we partition the molecule into a BPEGraph at a specific scale and serialize it via scaffold-rooted breadth-first search (BFS)  (Bundy and Wallen, 1984). This produces a CamS subsequence with a stable Intra-scale Order, moving from the global core to peripheral functional groups. Third, and most critically, to construct a hierarchical context, we concatenate subsequences from fine to coarse. This strategy circumvents the trade-off between scales and establishes an Inter-scale Order in the final CamS sequence, enabling high-level motifs to be predicted by conditioning on dense, uncorrupted fine-scale evidence. Collectively, these steps unify substructure compression and serialization through a dual-causal ordering—sequencing motifs within each scale and stacking scales hierarchically—to enable the model to comprehend global scaffolds based on explicit, uncorrupted local structural evidence.
@@
Built on CamS, we instantiate CamS-LLaMA by pre-training a native decoder-only backbone with standard NTP on CamS sequences, keeping the architecture and objective unchanged (Touvron et al., 2023). To keep attribution controlled, we inject only a lightweight molecular fingerprint (Cereto-Massagué et al., 2015) prior during downstream fine-tuning, while keeping pre-training purely sequence/structure-driven. Across MoleculeNet (Wu et al., 2018) and the activity-cliff stress test MoleculeACE (Van Tilborg et al., 2022), CamS-LLaMA achieves state-of-the-art performance at a comparable model scale to strong graph self-supervised learning baselines, despite relying on substantially weaker priors (Li et al., 2023). Beyond accuracy, we provide mechanism-level evidence. Targeted ablation studies confirm that multi-scale concatenation is the key driver of performance, while interpretability analysis reveals that the causal serialization effectively drives attention toward cliff-determining structural differences.
@@ 2.1 Molecular Property Prediction within Foundation Models Paradigm
Molecular property prediction is increasingly adopting a foundation model approach: large-scale Transformer pre-training followed by fine-tuning for diverse property endpoints (Awais et al., 2025; Xia et al., 2025). One approach adopts standard large language model (LLM) methodology by representing molecules as SMILES strings and pre-training decoder-only models via NTP (Xia et al., 2025; Cai et al., 2025). This leverages mature LLM infrastructure and enables straightforward large-scale training. However, SMILES is not structure-native, weakening topology-based signals. String-based edits often misalign with actual chemical changes, and performance consistently trails strong graph-native methods on property prediction tasks (Xia et al., 2025). Alternatively, graph-native foundation models directly process molecular graphs to preserve connectivity and local topology (You et al., 2021; Liu et al., ; You et al., 2020; Xu et al., 2021). High-performing methods typically employ encoder-only Graph Transformers (Shehzad et al., 2024) designed for property prediction. While effective, these models deviate from the vanilla decoder-only NTP approach that exhibits superior generalization and scalability (Li et al., 2023; Rong et al., 2020). More recent structural foundation models integrate graphs with autoregressive components but frequently depend on additional objectives (often incorporating 3D structure) and learned molecular encoders rather than pure tokenizer-level NTP (Kong et al., 2025; Lu et al., 2025; Zhang et al., 2025).
@@ 2.2 Molecular Tokenization Strategies
Within a foundation model paradigm, tokenizer design becomes a key interface for scaling LLMs. In NLP, BPE yields strong compression via simple frequency-based merges without training an extra tokenizer model. Molecular tokenization spans string-based and graph-based methods (Wang et al., 2019; Liu et al., 2023). String-based schemes operate on SMILES (Weininger, 1988) or related linearizations (Krenn et al., 2022), while graph-based schemes have evolved from basic node/edge-level units to fragment-level motifs (Liu et al., 2023). Some rule-based motif tokenizations rely on limited handcrafted chemistry rules (Zhang et al., 2021); triplet tokenizations, as used in KPGT (Knowledge-guided Pre-training of Graph Transformer), show that enriching a node token with only one extra atom can substantially boost property prediction (Li et al., 2023); ring-or-path-based (Wollschläger et al., 2024) and tree-based (Jin et al., 2018) schemes also expose substructures at different granularities; some other works learn tokenization via trainable rules (Sun et al., 2024) or by training a GNN as a tokenizer (Liu et al., 2023; Luo et al., ). Particularly noteworthy are data-driven motif mining approaches that adapt BPE idea to graphs to obtain reusable, multi-scale fragments with strong compression and minimal reliance on manually designed rules or additional trained models (Geng et al., 2023; Shen and Poczos, ; Kong et al., 2022). In addition, cross-level tokenization schemes, which combine motifs at different structural levels, also provide complementary design ideas (Ji et al., 2022; Chen et al., ).
@@ 3 Method
Our molecular representation must satisfy two requirements: (1) chemical fidelity, i.e., sensitivity to pharmacophore-level structures and their topology; and (2) autoregressive foundation model compatibility, i.e., a compressed interface naturally aligned with an NTP paradigm (Touvron et al., 2023). Motivated by this, we propose Connection-Aware Motif Sequencing (CamS), which segments molecular graphs into reusable frequent motifs and, crucially, serializes the motif graph into multi-scale connection-aware causal sequences compatible with NTP pre-training (Section 3.1). Built on CamS, CamS-LLaMA performs large-scale autoregressive pre-training and injects a only lightweight fingerprint prior during downstream fine-tuning (Section 3.2). Finally, we analyze the characteristics and advantages of multi-scale CamS autoregressive modeling by comparing it with Graphformer-style modeling (Section 3.3).
@@ 3.1.1 Substructure Level: CamS Vocabulary
BPE-style Mining. Following (Geng et al., 2023; Shen and Poczos, ), we first learn an ordered list of merge operations $\mathcal{O}$ from co-occurrence statistics. We then apply $\mathcal{O}$ to the corpus to mine the resulting subgraphs. This yields the BPEMotif set (frequent subgraphs with connectivity patterns) and the base BPE-SAV (Single-Atom Vocabulary with connectivity patterns) naturally captured during mining.
@@ 3.1.2 Per-Scale Encoding
Serialization (Intra-scale Order). To adapt the graph for decoder-only NTP, we serialize $V_{s}$ into an ordered list via Breadth-First Search (BFS). We select the motif with the largest atom count (typically the core scaffold) as the root. This strategy prioritizes global backbone structure before local substitutions (Zhang et al., 2021), establishing a stable Center-to-Periphery causal order.
@@ 3.2.1 Autoregressive Pre-training via NTP
CamS-LLaMA adopts a standard LLaMA-style decoder (Touvron et al., 2023); in pre-training, every CamS token list is treated as an ordinary token sequence in the shared vocabulary $\Sigma$ and is consumed by the backbone under the usual causal mask, with no graph-specific architectural modification. Concretely, each molecule yields five CamS sequence views $\{X_{1},X_{7},X_{27},X_{67},X\}$: four single-scale BFS sequences tokenized with $\Sigma_{1\mathrm{K}}/\Sigma_{7\mathrm{K}}/\Sigma_{27\mathrm{K}}/\Sigma_{67\mathrm{K}}$, plus the concatenated multi-scale sequence $X$ in Eq. (1); we treat each view as one NTP training instance (details in Appendix A.3). Given any view $X=(x_{1},\ldots,x_{T})$, token embeddings form $\mathbf{H}^{(0)}$, which is transformed by $L$ causal self-attention blocks into $\mathbf{H}^{(L)}$; a tied output projection then yields $p_{\theta}(x_{t+1}\mid x_{\leq t})$.
@@ 3.2.2 Fine-tuning with Fingerprint Injection
Many molecular pre-training methods inject rich handcrafted descriptors early (e.g., during pre-training) (Li et al., 2023). In contrast, we keep CamS as the primary representation and inject only one widely-used structural prior–a fingerprint–only during fine-tuning, to make the contribution of CamS representations more attributable and controlled.
@@
Graphformer (Flat/Static Bias). $\mathcal{V}_{\text{tok}}$ corresponds to single-scale atoms or triplets. $\psi_{ij}$ encodes static structural biases (e.g., shortest-path-distance bias injected into attention) independent of token content, yielding an isotropic receptive field at a given graph distance (Ying et al., 2021).
@@
Sketch. The inequality follows from the Data Processing Inequality (DPI): masking is a stochastic post-processing of $Z_{t}$ that cannot increase mutual information (Cover, 1999). In graph MNP, this loss of information is exacerbated by evidence-pattern uncertainty: the most relevant local-neighborhood evidence for a masked token may be co-masked, making the conditional distribution unstable across masking patterns. Recent graph pretraining works explicitly mitigate this by masking structured units (e.g., triplets / local subgraphs) or designing structure-guided masking, and by injecting global knowledge tokens (Li et al., 2023; Rong et al., 2020; Liu et al., 2024). In contrast, CamS NTP for coarse-scale tokens conditions on an uncorrupted fine-to-coarse history, avoiding such stochastic evidence loss. This perspective is consistent with classic critiques of Masked Language Modelling (MLM): due to input corruption and factorized prediction over masked positions, it introduces a pre-train–fine-tune discrepancy and neglects dependencies among masked targets (Yang et al., 2019).
@@
This trade-off (more predicted tokens vs. more corruption) is central to masked modeling (Wettig et al., 2023). Empirically, KPGT finds the best masking ratio can be as high as $\rho=0.5$ when using triplet-style tokens and global knowledge nodes to stabilize context (Li et al., 2023). Meanwhile, dense token-level learning signals are known to substantially improve sample efficiency compared to sparse masked-only supervision (Clark et al., 2020). CamS further strengthens this effect by providing multi-view supervision: the same molecule contributes token-level targets repeatedly across scales.
@@ 4.1 Experimental Setup
CamS-Tokenizer Training. CamS-tokenizer training is expensive because replaying $\mathcal{O}$ to materialize $\mathcal{V}$ repeatedly invokes RDKit’s subgraph-isomorphism matching (the dominant bottleneck) (Landrum et al., 2021; Ehrlich and Rarey, 2011) to extract motifs and attachment patterns, which is difficult to accelerate. We therefore train the tokenizer once on ChEMBL-34 (Gaulton et al., 2017), a moderately sized yet chemically diverse corpus, at the largest scale (67K). Concretely, we mine an ordered merge list $\mathcal{O}=(o_{1},\dots,o_{K})$ with K=685K{=}685 merges, apply it to obtain $\mathcal{V}$, and then perform SAVC to ensure single-atom coverage for common elements. Smaller scales are obtained without retraining by prefix slicing: each scale uses a prefix $\mathcal{O}_{\leq k}$ together with its induced vocabulary subset (i.e., motifs mined by applying $\mathcal{O}_{\leq k}$), where 1K uses k=0k{=}0, 7K/27K use k=62/210k{=}62/210, and 67K uses k=685k{=}685. Implementation details and full procedures are provided in Appendix A.1.
Pre-training Dataset and Configuration. We instantiate a 16-layer LLaMA-style decoder (hidden size 720, 8 attention heads) using HuggingFace Transformers. For self-supervised pre-training, we use Enamine675M (675M molecules) (Shivanyuk et al., 2007) and optimize NTP (Section 3.2.1). For each molecule, we construct five CamS training sequences: four single-scale BFS sequences tokenized with 1K/7K/27K/67K vocabularies, plus the concatenated multi-scale sequence in Eq. (1). We treat each sequence as one training instance and uniformly shuffle all instances, resulting in roughly $5\times$ as many sequences as molecules. All optimization details (schedule, batch size, steps) are summarized in Appendix C.1.
Downstream Benchmarks and Baselines. We evaluate on MoleculeNet (11 tasks; broad endpoints that probe general-purpose molecular property prediction) (Wu et al., 2018) and MoleculeACE (30 activity-cliff tasks; stress-testing fine-grained structural sensitivity under subtle edits that induce large activity shifts) (Van Tilborg et al., 2022). To enable an apples-to-apples comparison against a recent strong pre-training baseline, we adopt the data splits and evaluation protocol used in KPGT (Li et al., 2023). On MoleculeNet, following KPGT’s setup, we compare primarily against self-supervised pre-training baselines to isolate the effect of (1) representation/tokenization and (2) the pre-training objective. On MoleculeACE, we report the benchmark’s comprehensive ML/DL baselines, since activity cliffs emphasize discriminating small structural changes where strong feature-based methods remain competitive. Unless stated otherwise, CamS-LLaMA is kept at $\sim$100M parameters (comparable to KPGT); fingerprints are injected only during fine-tuning as a minimal prior, while pre-training is purely sequence-based.